# Minghan Chen  Guikun Chen  Wenguan Wang  Yi Yang

### v1

@@ 4.1 Experimental Setup
  • Visual Genome (VG150) [28; 68] contains 150 entity and 50 relation categories. It is split into 57,723 training, 5,000 validation, and 26,446 testing images.
@@
Competitors. Hydra-SGG is compared with methods from two categories. The first category contains two-stage SGG methods, including MOTIFS [77], VCTree-TDE [60], BGNN [36], PE-Net [83], IS-GGT [30], VETO [56], and UniVRD [82], CFA [35], SHA [8]. The second consists of one-stage methods including SGTR [37], ISG [25], RelTR [6], DSGG [12], SpeaQ [27], EGTR [18].
Implementation Details. Following the commonly used settings of DETR-like SGG models [37; 6], we employ ResNet50 [13] as our backbone. We implement various DETR-based techniques in our baseline and proposed models, including anchor boxes, training techniques, and attention mechanisms as used in recent works [6; 45; 32; 81]. Our model utilizes the same One-to-One Relation Assignment strategy from RelTR [6] and matching hyperparameters. Our model architecture consists of a 6-layer Transformer encoder and a 6-layer Transformer decoder, each with 256 dimensions and 8 attention heads, which is the same as previous studies [6; 81; 45]. The prediction heads are shared for RelDecoder and HydraBranch. All models are trained from scratch using AdamW optimizer with an initial learning rate of 2e-4 and the total batch size is 16. More implementation details, including hyperparameter settings and data preprocessing, are provided in Appendix.
@@ 4.2 Quantitative Comparison Result
VG150 [68] test. Table 1 reports the comparison results on VG150 test. Hydra-SGG demonstrates outstanding performance on challenging SGDet, achieving mR@20 and mR@50 scores of 10.6 and 16.0, respectively, setting a new state-of-the-art. Hydra-SGG achieves this performance with significantly shorter training time, requiring only 12 epochs. This training time is substantially shorter compared to SpeaQ [27], EGTR [18], and DSGG [12], with reductions of 40, 263, and 48 epochs, respectively. Despite the shorter training, Hydra-SGG outperforms these methods by +4.2, +8.1, and +3.0 on the mR@50. Furthermore, it even outperforms the state-of-the-art two-stage model, PE-Net [83], by a margin of 3.6 on the mR@50 metric. It is crucial to acknowledge that the backbone model employed by DSGG [12] is not disclosed in their paper, and the authors could not provide this information upon request. This raises concerns about the fairness of the comparison, as DSGG [12] may use Swin-L [46] as the backbone model.
Table 1: SGDet evaluation on VG150 [28] test4.2). The methods are categorized into two- and one-stage groups. Epochs marked with + include detector pre-training time on VG150. FPS (Frames Per Second) indicates inference speed. Higher FPS values represent faster inference.
       Method Backbone # Epoch FPS # Param (M) R@20 R@50 R@100 mR@20 mR@50 mR@100
@@
MOTIFS [77] [CVPR2018] ResNeXt101-FPN - - 369.9 25.1 32.1 36.9 4.1 5.5 6.8
VCTree-TDE [60] [CVPR2020] ResNeXt101-FPN - - 361.3 14.3 19.6 - 6.3 9.3 11.1
BGNN [36] [CVPR2021] ResNeXt101-FPN - - 341.9 23.3 31.0 35.8 7.5 10.7 12.7
PE-Net [83] [CVPR2023] ResNeXt101-FPN 32+ - - - 30.7 35.2 - 12.4 14.5
IS-GGT [30] [CVPR2023] ResNet101 70 - - - - - - 9.1 11.3
VETO [56] [ICCV2023] ResNeXt101-FPN 33+ - - - 27.5 31.5 - 8.1 9.5
UniVRD [82] [ICCV2023] CLIP ViT-B - - - - - - - 9.6 12.1
DRM [33] [CVPR2024] ResNeXt101-FPN - - - - 34.0 38.9 - 9.0 11.2
@@
SGTR [37] [CVPR2022] ResNet101 123+ - 117.1 - 25.1 26.6 - 12.0 14.6
ISG [25] [NeurIPS2022] ResNet101 52 - 93.5 - 29.5 32.1 - 7.4 8.4
RelTR [6] [TPAMI2023] ResNet50 150 6.5 63.7 21.2 27.5 30.7 6.8 10.8 12.3
DSGG [12] [CVPR2024] - 60 - - - 32.9 38.5 - 13.0 17.3
SpeaQ [27] [CVPR2024] ResNet101 52 - - - 32.9 36.0 - 11.8 14.1
EGTR [18] [CVPR2024] ResNet50 275+ - 42.5 23.5 30.2 34.3 5.5 7.9 10.1
@@
Hydra-SGG   ResNet50 12 5.3 67.6 21.9 28.4 33.1 10.6 16.0 19.7
[…]
@@
Motifts [77] [CVPR2018] ResNeXt101-FPN - 369.9 71.6 29.9 31.6 38.9
BGNN [36] [CVPR2021] ResNeXt101-FPN - 341.9 75.0 35.5 34.2 42.1
@@
SGTR [37] [CVPR2022] ResNet101 123+ 117.1 59.9 37.0 38.7 42.3
RelTR [6] [TPAMI2023] ResNet50 150 63.7 71.7 34.2 37.5 43.0
EGTR [18] [CVPR2024] ResNet50 275+ 42.5 75.0 42.0 41.9 48.6
@@
Hydra-SGG   ResNet50 7 67.6 76.1 42.8 44.3 50.1
@@
     Method Backbone R@50 R@100 mR@50 mR@100
SHA [8] [CVPR2022] ResNeXt101-FPN 25.5 29.1 6.6 7.8
VETO [56] [ICCV2023] ResNeXt101-FPN 26.1 29.0 7.0 8.1
CFA [35] [ICCV2023] ResNeXt101-FPN - - 10.8 12.6
@@
Hydra-SGG   ResNet50 22.8 26.5 12.7 15.9
GQA [16] test. Our experimental results demonstrate that Hydra-SGG achieves SOTA mean Recall performance on GQA. As shown in Table 3, HydraSGG achieves 12.7 and 15.9 for mR@50 and mR@100 respectively, surpassing previous best results. Our model employs the relatively lightweight ResNet50 as its backbone, in contrast to the heavier ResNeXt101 used by other methods. Notably, CFA [35] achieves good mR results using unbiasing techniques, but Hydra-SGG surpasses these without such methods. Our approach outperforms CFA by 1.9 and 3.3 in mR@50 and mR@100 respectively, demonstrating its effectiveness.
@@ 4.3 Diagnostic Experiment
We will conduct ablation studies to address the following key questions: (1) How effective are the core components of Hydra-SGG? (2) How does the training efficiency of Hydra-SGG compare to existing methods? (3) What is the impact of threshold $T$ in One-to-Many Relation Assignment on model performance? (4) How do the number of training epochs influence model performance? (5) What is the effect of the number of relation queries on model performance? Through these experiments, we aim to comprehensively evaluate various aspects of Hydra-SGG.
Key Components. We first investigate the effectiveness of our core Hybrid Relation Assignment (§3.2) and Hydra Branch (§3.3) in Table 4(a). The first row gives the score of our baseline model (§3.1). The second row corresponds to the result of Vanilla Hydra-SGG, which directly utilizes Hybrid Relation Assignment into RelDecoder. The third row lists the results of the full Hydra-SGG model. We can find that taking the proposed Hybrid Relation Assignment in the baseline model improves the mean recall and achieves 9.9/16.0 mR@20/mR@50. Furthermore, the integration of Hydra Branch with One-to-Many Relation Assignment yields a synergistic effect, boosting performance substantially. Specifically, Hydra Branch increases mR@20 from 9.9 to 10.6 and mR@50 from 14.9 to 16.0, highlighting the complementary nature of these two strategies.
Training Efficacy. In Table 4(b), we compare the training time cost on VG150 [68]. All models are trained on eight NVIDIA 4090 GPUs with a ResNet50 backbone. Our model achieves 16.0 mR@50 in 720 minutes, whereas RelTR takes 3,000 minutes to achieve 6.8 mR@50. This demonstrates the improved training efficacy of Hydra-SGG.
@@ 4.4 Qualitative Comparison Result
In ${}_{\!}$ 5a-d, we visualize the generated scene graph results on the image of VG150 [28] val. Both Hydra-SGG and RelTR [6] detect plausible relations, but these relations were not annotated, so these reasonable predictions become false negatives. Our Hydra-SGG detects more fine-grained relations such as rock-on-mountain and logo-on-shirt, while RelTR fails to detect such relations. In ${}_{\!}$ 5e-g, we visualize generated scene graph results on Unsplash images. Hydra-SGG successfully identifies a new species of bird that was not present in the training dataset in ${}_{\!}$ 5e. In ${}_{\!}$ 5f, the model recognizes a truck and a person standing on it against a starry sky background, even though only the black silhouettes of the person and truck are visible. In ${}_{\!}$ 5g, the model successfully identifies trees on a distant mountain, as well as tree-behind-man. This demonstrates the generalization of our method to unseen web images.
@@ 5 Conclusion
This paper introduces Hydra-SGG, a one-stage DETR-based scene graph generation (SGG) model that addresses the slow convergence issue in existing DETR-based SGG models. The key contributions of our work are as follows: i) We propose a Hybrid Relation Assignment, which combines One-to-One and One-to-Many Relation Assignment strategies to increase relation supervision signals. ii) We introduce a Hydra Branch, an auxiliary decoder that encourages relation queries to predict duplicate relations, further enhancing the proposed One-to-Many Relation Assignment. These innovations work in synergy to significantly accelerate the learning process. Consequently, Hydra-SGG achieves state-of-the-art results on VG150, GQA, and Open Images V6 in just 12, 12, and 7 training epochs, demonstrating remarkable performance and efficiency. Despite these advancements, there are still opportunities for further improvement. Future work should explore integrating Large Language Models (LLMs) with Hydra-SGG to enable open vocabulary SGG. This combination could potentially allow the model to recognize and describe a much wider range of objects and relationships beyond predefined categories.
Appendix A Appendix
@@
  • §B details the implementation details.
  • §C provides the pseudo code of Hydra-SGG.
  • §D shows failure cases of Hydra-SGG in VG150 [68].
  • §E discusses our limitations, societal impact, and directions of future work.
Appendix B Implementation
Inspired by RelTR [6], we also set relation queries to be composed of subject and object queries. We adopt the same training techniques and attention mechanisms as previous works [32; 81; 6] in Hydra-SGG. In particular, we use anchor boxes as embeddings to accelerate training and employ deformable attention [84] respectively. We input noised G) box-label pairs to generate both entity and relation denoising queries. To prevent information leakage, we employ attention masks. Given the absence of an attention map in the deformable decoder, we simplify the model by omitting the convolutional mask head in RelTR [6]. In particular, we concatenate the output subject and object queries and input them into an MLP. This simplification not only reduces computational complexity but also maintains performance.
Both One-to-One and One-to-Many losses are composed of box regression loss, entity classification loss, and relation classification loss as previous works [2; 37; 6; 18; 58]. Specifically, we use Focal loss [42] for both relation and detection, as in previous works [45; 32; 81]. In the post-processing stage, the predictions are first ranked by the relation probability. Then we use the relation index to find corresponding subject and object predictions to compose the prediction triplets. Finally, we follow the same process as RelTR [6] that removes predictions where the subject and object are the same, as such data does not exist in VG150 [68], but we do not use this process in Open Images V6 [31].
@@ Appendix B Implementation
Due to the inherent bias in SGG datasets, numerous unbiasing algorithms have been developed [60; 35; 34; 56; 83]. While adopting these unbiasing techniques typically leads to significant performance improvements, we opt for a fair comparison by evaluating Hydra-SGG against methods that do not employ such techniques.
Appendix C Pseudo Code
The pseudo-code of Hydra-SGG is given in Algorithm 1. Our code and pre-trained models will be made publicly available.

### v2

@@ 4.1 Experimental Setup
  • Visual Genome (VG150) [27; 75] contains 150 entity and 50 relation categories. It is split into 57,723 training, 5,000 validation, and 26,446 testing images.
@@
Competitors. Hydra-SGG is compared with methods from two categories: (1) Two-stage SGG methods, including MOTIFS [85], VCTree-TDE [65], BGNN [39], PE-Net [91], IS-GGT [29], VETO [61], UniVRD [90], DRM [32], CFA [34], and SHA [9]; (2) One-stage methods including SGTR [40], SSR-CNN [67], ISG [24], RelTR [7], DSGG [13], SpeaQ [26], and EGTR [18].
@@ 4.2 Quantitative Comparison Result
VG150 [75] test. Table 1 reports the comparison results on VG150 test. Hydra-SGG demonstrates outstanding performance on challenging SGDet, achieving mR@20 and mR@50 scores of 10.6 and 16.0, respectively, setting a new state-of-the-art. Hydra-SGG achieves this performance with significantly shorter training time, requiring only 12 epochs. This training time is substantially shorter compared to SpeaQ [26], EGTR [18], and DSGG [13], with reductions of 40, 263, and 48 epochs, respectively. Despite the shorter training, Hydra-SGG outperforms these methods by +4.2, +8.1, and +3.0 on the mR@50. Furthermore, it even outperforms the state-of-the-art two-stage model, PE-Net [91], by a margin of 3.6 on the mR@50 metric. We speculate that the significant improvement in mR can be attributed to our One-to-Many Assignment strategy, which ensures balanced supervision across both rare and common relations by allocating a fixed number of six queries per relation category (§ 3.2). For instance, in an image containing ten “on” relations and two “sit” relations, while One-to-One Assignment would allocate ten queries to “on” and only two to “sit”, our approach assigns six queries to each, resulting in a 60% increase for common relations and a substantial 300% increase for rare relations.
Table 1: SGDet evaluation on VG150 [27] test4.2). +: detector pre-trained on VG150. FPS (Frames Per Second) indicates inference speed. F-Recall of Hydra-SGG is calculated based on the best results.
       Method Backbone # Epoch FPS # Param R@20/50/100 mR@20/50/100 F@20/50/100
@@
MOTIFS [85] [CVPR2018] ResNeXt101-FPN - - 369.9M 25.1 / 32.1 / 36.9 4.1 / 5.5 / 6.8 7.1 / 9.2 / 11.7
VCTree-TDE [65] [CVPR2020] ResNeXt101-FPN - - 361.3M 14.3 / 19.6 / - 6.3 / 9.3 / 11.1 8.8 / 12.4 / -
BGNN [39] [CVPR2021] ResNeXt101-FPN - - 341.9M 23.3 / 31.0 / 35.8 7.5 / 10.7 / 12.7 11.3 / 15.5 / 19.0
PE-Net [91] [CVPR2023] ResNeXt101-FPN 32+ - - - / 30.7 / 35.2 - / 12.4 / 14.5 - / 17.7 / 21.2
IS-GGT [29] [CVPR2023] ResNet101 70 - - - / - / - - / 9.1 / 11.3 - / - / -
VETO [61] [ICCV2023] ResNeXt101-FPN 33+ - - - / 27.5 / 31.5 - / 8.1 / 9.5 - / 12.5 / 14.6
UniVRD [90] [ICCV2023] CLIP ViT-B - - - - / - / - - / 9.6 / 12.1 - / - / -
DRM [32] [CVPR2024] ResNeXt101-FPN - - - - / 34.0 / 38.9 - / 9.0 / 11.2 - / 14.2 / 17.4
@@
SGTR [40] [CVPR2022] ResNet101 123+ - 117.1M - / 25.1 / 26.6 - / 12.0 / 14.6 - / 16.2 / 18.9
SSR-CNN [67] [CVPR2022] ResNet101 - - 274.3M 25.8 / 32.7 / 36.9 6.1 / 8.4 / 10.0 9.9 / 13.4 / 15.7
ISG [24] [NeurIPS2022] ResNet101 52 - 93.5M - / 29.5 / 32.1 - / 7.4 / 8.4 - / 11.8 / 13.3
RelTR [7] [TPAMI2023] ResNet50 150 6.5 63.7M 21.2 / 27.5 / 30.7 6.8 / 10.8 / 12.3 10.3 / 15.5 / 17.6
DSGG [13] [CVPR2024] - 60 - - - / 32.9 / 38.5 - / 13.0 / 17.3 - / 18.6 / 23.9
SpeaQ [26] [CVPR2024] ResNet101 52 - - - / 32.9 / 36.0 - / 11.8 / 14.1 - / 17.4 / 20.3
EGTR [18] [CVPR2024] ResNet50 275+ 7.7 42.5M 23.5 / 30.2 / 34.3 5.5 / 7.9 / 10.1 8.9 / 12.5 / 15.6
@@
Hydra-SGG  [ICLR2025] ResNet50 12 5.3 67.6M 21.9 / 28.6 / 33.4 $\pm$0.1 / $\pm$0.2 / $\pm$0.3 10.3 / 15.9 / 19.4 $\pm$0.2 / $\pm$0.2 / $\pm$0.2 14.0 / 20.5 / 24.7
Table 2: Evaluation on Open Images V6 [30] test4.2). +: detector pre-trained on Open Images V6.
@@
Motifts [85] [CVPR2018] ResNeXt101-FPN - 369.9M 71.6 29.9 31.6 38.9
BGNN [39] [CVPR2021] ResNeXt101-FPN - 341.9M 75.0 35.5 34.2 42.1
@@
SGTR [40] [CVPR2022] ResNet101 123+ 117.1M 59.9 37.0 38.7 42.3
RelTR [7] [TPAMI2023] ResNet50 150 63.7M 71.7 34.2 37.5 43.0
EGTR [18] [CVPR2024] ResNet50 275+ 42.5M 75.0 42.0 41.9 48.6
@@
Hydra-SGG  [ICLR2025] ResNet50 7 67.6M $76.0_{\pm 0.2}$ $\textbf{42.8}_{\pm 0.2}$ $\textbf{44.1}_{\pm 0.2}$ $\textbf{50.0}_{\pm 0.2}$
@@
     Method Backbone R@50/100 mR@50/100
SHA [9] [CVPR2022] ResNeXt101-FPN 25.5 / 29.1 6.6 / 7.8
VETO [61] [ICCV2023] ResNeXt101-FPN 26.1 / 29.0 7.0 / 8.1
CFA [34] [ICCV2023] ResNeXt101-FPN - 10.8 / 12.6
@@
Hydra-SGG  [ICLR2025] ResNet50 23.1 / 26.8 $\pm$0.3 / $\pm$0.3 12.5 / 15.6 $\pm$0.2 / $\pm$0.3
GQA [16] test. Our experimental results demonstrate that Hydra-SGG achieves SOTA mean Recall performance on GQA. As shown in Table 3, Hydra-SGG achieves 12.7 and 15.9 for mR@50 and mR@100 respectively, surpassing previous best results. Our model employs the relatively lightweight ResNet50 as its backbone, in contrast to the heavier ResNeXt101 used by other methods.
Comparisons with Unbiasing Methods. As shown in Table 4(g), Hydra-SGG achieves very competitive performance on VG150 compared to these specialized debiasing methods, including TDE [65], NICE [33], IETrans [87], CFA [34], VETO [61], and NICEST [36].
@@
       Loss Ratio (1-to-1 : 1-to-M) mR@50 mR@100
1 : 0.5 14.6 18.9
1 : 0.8 15.6 19.4
1 : 1 16.0 19.7
1 : 1.5 15.4 18.7
1 : 2 15.6 19.5
1 : 3 15.4 19.3
(f) Loss Weights Between One-to-One (1-to-1)
and One-to-Many (1-to-M) Losses.
          Method mR@50 mR@100 TDE [65] [CVPR2020] 9.2 11.1 NICE [33] [CVPR2022] 10.4 12.7 IETrans [87] [ECCV2022] 12.5 15.0 CFA [34] [ICCV2023] 12.3 14.6 VETO [61] [ICCV2023] 10.6 13.8 NICEST [36] [TPAMI2024] 10.4 12.4 Hydra-SGG [ICLR2025] 16.0 19.7
(g) Comparisons with Unbiasing Methods.
@@ 4.3 Diagnostic Experiment
Key Components. We first investigate the effectiveness of our core Hybrid Relation Assignment (§3.2) and Hydra Branch (§3.3) in Table 4(a). The first row gives the score of our baseline model (§3.1). The second row corresponds to the result of Vanilla Hydra-SGG, which directly applies Hybrid Relation Assignment into RelDecoder. The third row lists the results of the complete Hydra-SGG model. Our results show that taking the proposed Hybrid Relation Assignment in the baseline model improves the mean recall and achieves 9.9/14.9 mR@20/mR@50. Furthermore, the integration of Hydra Branch with Hybrid Relation Assignment yields a synergistic effect, boosting performance substantially. Specifically, Hydra Branch increases mR@20 from 9.9 to 10.6 and mR@50 from 14.9 to 16.0, highlighting the complementary nature of these two strategies.
Training Efficacy. In Table 4(b), we compare the training time costs on VG150 [75]. All models are trained on eight NVIDIA RTX 4090 GPUs with a ResNet50 backbone. Our model achieves 16.0 mR@50 in 12 hours, whereas RelTR takes 50 hours to achieve 6.8 mR@50. This demonstrates the superior training efficacy of Hydra-SGG.
@@
Impact of Loss Weights. As shown in Table 4(f), we conducted an ablation study to investigate the effect of different loss weights between One-to-One and One-to-Many losses. With a 1:1 loss ratio, the model achieves the best results (mR@50: 16.0, mR@100: 19.7). When the One-to-Many loss weight is reduced (e.g., 1:0.5), the model fails to fully utilize the additional positive samples available. Conversely, increasing this weight excessively (e.g., 1:3) leads to an overemphasis on One-to-Many triplets, which typically have lower quality compared to One-to-One triplets. This quality difference arises from the fundamental nature of the assignments: One-to-One assignment matches each relation label with its optimal query exclusively, while One-to-Many assignment matches a relation label with multiple queries meeting certain criteria.
@@ 4.4 Qualitative Comparison Result
In ${}_{\!}$ 5a-d, we visualize the generated scene graph results on the image of VG150 [27] val. Both Hydra-SGG and RelTR [7] detect plausible relations, but these relations were not annotated, so these reasonable predictions become false negatives. Our Hydra-SGG detects more fine-grained relations such as rock-on-mountain and logo-on-shirt, while RelTR fails to detect such relations. In ${}_{\!}$ 5e-g, we visualize generated scene graph results on unseen Unsplash images.
@@ 5 Conclusion
This paper introduces Hydra-SGG, a one-stage DETR-based scene graph generation (SGG) model that addresses the slow convergence issue in existing DETR-based SGG models. The key contributions of our work are as follows: i) We propose a Hybrid Relation Assignment, which combines One-to-One and One-to-Many Relation Assignment strategies to increase relation supervision signals. ii) We introduce a Hydra Branch, an auxiliary decoder that encourages relation queries to predict duplicate relations, further enhancing the proposed One-to-Many Relation Assignment. These innovations work in synergy to significantly accelerate the learning process. Consequently, Hydra-SGG achieves state-of-the-art results on VG150 [75], GQA [16], and Open Images V6 [30] in just 12, 12, and 7 training epochs, demonstrating remarkable performance and efficiency.
Appendix
@@
  • §A details the implementation details.
  • §B provides the pseudo code of Hydra-SGG.
  • §C shows failure cases of Hydra-SGG in VG150 [75].
  • §D discusses our limitations, societal impact, and directions of future work.
Appendix A Implementation
Inspired by RelTR [7], we also set relation queries to be composed of subject and object queries. We adopt the same training techniques and attention mechanisms as previous works [31; 89; 7] in Hydra-SGG. In particular, we use anchor boxes as embeddings to accelerate training and employ deformable attention [93] respectively. We input noised box-label pairs to generate both entity and relation denoising queries. To prevent information leakage, we employ attention masks. Given the absence of an attention map in the deformable decoder, we simplify the model by omitting the convolutional mask head in RelTR [7]. In particular, we concatenate the output subject and object queries and input them into an MLP. This simplification not only reduces computational complexity but also maintains performance.
Both One-to-One and One-to-Many losses are composed of box regression loss, entity classification loss, and relation classification loss as previous works [2; 40; 7; 18; 63]. Specifically, we use Focal loss [45] for both relation and detection, as in previous works [49; 31; 89]. In the post-processing stage, the predictions are first ranked by the relation probability. Then we use the relation index to find corresponding subject and object predictions to compose the prediction triplets. Finally, we follow the same process as RelTR [7] that removes predictions where the subject and object are the same, as such data does not exist in VG150 [75], but we do not use this process in Open Images V6 [30].
@@ Appendix A Implementation
Due to the inherent bias in SGG datasets, numerous unbiasing algorithms have been developed [65; 34; 33; 61; 91]. While adopting these unbiasing techniques typically leads to significant performance improvements, we opt for a fair comparison by evaluating Hydra-SGG against methods that do not employ such techniques.
Appendix B Pseudo Code
The pseudo-code of Hydra-SGG is given in Algorithm S1. Our code and pre-trained models will be made publicly available.